# Learning About Structural Errors in Models of Complex Dynamical Systems

### v1

@@ 4.2.2 Lorenz 96 Single-scale Model
We studied the Lorenz 96 single-scale system to learn the quadratic term as an error model. Using this numerical example, we demonstrate the merit of combined use of direct and indirect data and the importance of enforcing physical constraints. More specifically, we assume that the quadratic term of the true system in (25) is unknown and then calibrate an error model $\delta(X_{k-2},X_{k-1},X_{k+1},X_{k+2})$ as in (26). The numerical examples for the single-scale Lorenz 96 system are as follows:
@@ 5.1 Ultradian Model
We consider the ultradian model of the glucose-insulin system proposed in Sturis et al. (1991). Its primary state variables are the plasma glucose concentration, $G$, the plasma insulin concentration, $I_{p}$, and the interstitial insulin concentration, $I_{i}$. These three state variables are augmented with a three stage delay $(h_{1},h_{2},h_{3})$ which encodes a non-linear delayed hepatic glucose response to plasma insulin levels. The resulting ordinary differential equations have the form:
@@
We represent a potential structural error in this model by removing the final linear term in (27a) (i.e., $t_{p}=\infty$). We then aim to identify a neural-network-based additive correction term for (27a), denoted by $\delta(I_{p},I_{i},G,h_{1},h_{2},h_{3};\theta)$. We calibrate $\delta(\cdot\ ;\ \theta)$ using partial, noisy observations of the physiologic state at times $\{t_{k}\}_{k=0}^{K}$. Specifically, we observe only the glucose state, $G(t)$, (it is the only state variable that is reliably measurable in patients) and add i.i.d. unit-Gaussian noise. Thus the observation operator $\mathcal{H}$ in Eq. (2) is defined through observation matrix $H\in\mathbb{R}^{1\times 6}$. To be specific $HX(t)=X_{3}(t)$ where $X(t)=[I_{p}(t),I_{i}(t),G(t),h_{1}(t),h_{2}(t),h_{3}(t)]^{T}$ and then $\mathcal{H}[X]=\{HX(t_{k})\}_{k=0}^{K}.$ It should be noted that the observation matrix $H$ projects between the full system state space and the observable space and may be used in the definition of a gain matrix $K$ for Kalman methods and generalizations such as 3DVAR (See Table 9.1 in Sanz-Alonso et al. (2023)).
@@
Specifically $\mathcal{G}_{k}(\theta;\ y)$ uses sequential data assimilation to combine previous observations $\big{[}y(t_{1}),\cdots,y(t_{k-1})\big{]}$ with parameters $\theta$ to produce a filtered estimate for $x(t_{k-1})$ and, from this, predict the observation $y(t_{k})$; see section 2.2.3. The minimization of Eq. (28) is performed using EKI, where each evaluation of $\mathcal{G}(\theta;y)$ involves running a data assimilation sub-routine (3DVAR Sanz-Alonso et al. (2023) with constant gain $K=[0,0,1,0,0,0]$).
@@ 5.2 Numerical Results for Ultradian Model
Figure 12 presents results from a trained model obtained by minimizing (28) via EKI. Figure 12c shows the transient trajectory data used to calibrate the model error, along with the obtained next-step predictions, which are seen to be highly accurate. Figure 12a shows that the neural network captures the true missing linear term $-I_{p}/t_{p}$ quite accurately in a wide range of inputs $I_{p}$, but is inaccurate where no data are available. Figure 12b shows the neural network as a function of an inactive input $G$ (an input upon which the true missing term does not depend); while it learns to correctly ignore $G$ in some of the plotted scenarios, it has spurious behavior outside of the observed range ($>200$) of $G$. Other results (not shown) indicate that the neural network learned substantial dependencies on inactive variables. This suggests that sparsity constraints may be useful in these experiments.
@@ Acknowledgments
TS, AMS, and JW were supported by the generosity of Eric and Wendy Schmidt by recommendation of the Schmidt Futures program. AMS is also supported by the Office of Naval Research (ONR) through grant N00014-17-1-2079 and by a Department of Defense Vannevar Bush Faculty Fellowship. MEL was supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1745301 and by funding from the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard.
@@ Appendix A Ensemble Kalman inversion
The use of ensemble Kalman based methods for parameter calibration and the solution of inverse problems, and history of this subject, is overviewed in [Section 4]Calvello et al. (2022). To be concrete we will concentrate on a particular variant of the methodology, sometimes termed Ensemble Kalman inversion (EKI). This is a specific ensemble-based, gradient-free optimization scheme that was proposed and studied in Iglesias et al. (2013); we emphasize that other ensemble Kalman based methods share the core desirable attributes of EKI, namely that it is derivative-free, is effective with relatively few evaluations of the forward model $\mathcal{G}$ and is robust to the presence of noise in the evaluations of $\mathcal{G}.$
The core task of EKI is equivalent to a quadratic optimization problem, which facilitates adding linear equality and inequality constraints Albers et al. (2019). To explain the details of EKI, we first introduce a new variable $w=\mathcal{G}(\theta)$ and variables $v$ and $g(v)$:
@@
to threshold those $\theta_{i}$ with values close to zero, after having solved the constrained optimization problem in (33). Such a thresholding step after the $\ell_{1}$-constrained optimization in (33) is equivalent to adding $\ell_{0}$ constraint. More details about imposing sparsity into EKI can be found in Schneider et al. (2022).
@@ Appendix B Application to a Glucose-Insulin Model
The functional forms of the parameterized processes in Eq. (5.1) are as follows:

### v2

@@ 4.2.2 Lorenz 96 Single-scale Model
We studied the Lorenz 96 single-scale system to learn the quadratic term as an error model. Using this numerical example, we demonstrate the merit of combined use of direct and indirect data and the importance of enforcing physical constraints. More specifically, we assume that the quadratic term of the true system in (25) is unknown and then calibrate an error model $\delta(X_{k-2},X_{k-1},X_{k+1},X_{k+2})$ as in (26). The size of the direct data is 36000, and the size of the indirect data is 154. The numerical examples for the single-scale Lorenz 96 system are as follows:
@@ 5.1 Ultradian Model
We consider the ultradian model of the glucose-insulin system proposed in [123]. Its primary state variables are the plasma glucose concentration, $G$, the plasma insulin concentration, $I_{p}$, and the interstitial insulin concentration, $I_{i}$. These three state variables are augmented with a three stage delay $(h_{1},h_{2},h_{3})$ which encodes a non-linear delayed hepatic glucose response to plasma insulin levels. The resulting ordinary differential equations have the form:
@@
We represent a potential structural error in this model by removing the final linear term in (27a) (i.e., $t_{p}=\infty$). We then aim to identify a neural-network-based additive correction term for (27a), denoted by $\delta(I_{p},I_{i},G,h_{1},h_{2},h_{3};\theta)$. We calibrate $\delta(\cdot\ ;\ \theta)$ using partial, noisy observations of the physiologic state at times $\{t_{k}\}_{k=0}^{K}$. Specifically, we observe only the glucose state, $G(t)$, (it is the only state variable that is reliably measurable in patients) and add i.i.d. unit-Gaussian noise. Thus the observation operator $\mathcal{H}$ in Eq. (2) is defined through observation matrix $H\in\mathbb{R}^{1\times 6}$. To be specific $HX(t)=X_{3}(t)$ where $X(t)=[I_{p}(t),I_{i}(t),G(t),h_{1}(t),h_{2}(t),h_{3}(t)]^{T}$ and then $\mathcal{H}[X]=\{HX(t_{k})\}_{k=0}^{K}.$ It should be noted that the observation matrix $H$ projects between the full system state space and the observable space and may be used in the definition of a gain matrix $K$ for Kalman methods and generalizations such as 3DVAR (See Table 9.1 in [124]).
@@
Specifically $\mathcal{G}_{k}(\theta;\ y)$ uses sequential data assimilation to combine previous observations $\big{[}y(t_{1}),\cdots,y(t_{k-1})\big{]}$ with parameters $\theta$ to produce a filtered estimate for $x(t_{k-1})$ and, from this, predict the observation $y(t_{k})$; see section 2.2.3. The minimization of Eq. (28) is performed using EKI, where each evaluation of $\mathcal{G}(\theta;y)$ involves running a data assimilation sub-routine (3DVAR [124] with constant gain $K=[0,0,1,0,0,0]$).
@@ 5.2 Numerical Results for Ultradian Model
  4. 4.
    Appropriate choice of regularization can improve the inference of otherwise un-identifiable model error structure (Figure 13).
Figure 12 presents results from a trained model obtained by minimizing (28) via EKI. Figure 12c shows the transient trajectory data used to calibrate the model error, along with the obtained next-step predictions, which are seen to be highly accurate. Figure 12a shows that the neural network captures the true missing linear term $-I_{p}/t_{p}$ quite accurately for $I_{p}\in[0,110]$, but it is inaccurate for larger values of $I_{p}$. While only $G$ is observed (not $I_{p}$), we note that indeed $I_{p}(t)<110$ for solutions of the true model. In other words, under the true model, the error term is only identifiable within the domain $[0,110]$. Figure 12b shows the neural network as a function of an inactive input $G$ (an input upon which the true missing term does not depend); instead, it learns a linear function of $G$ with a small slope. Other results (not shown) indicate that the neural network learned substantial dependencies on inactive variables. This suggests that sparsity constraints may be useful in these experiments.
@@
(a) Trained model: $\delta$ vs $I_{p}$
(b) Trained model: $\delta$ vs $G$
(c) Inferred parameter $\theta^{(\mathrm{NN})}_{20}$
Figure 13: Partial, noisy, transient timeseries data were used to calibrate model error with sparse EKI. Results in (a-b) show the trained neural network error model for two specific input scenarios and (c) shows an example of how sparse EKI drives initially non-zero error model coefficients to zero. Figures (a-b) plot the model $\delta$ (a function of all 6 state variables) for 5 fixed states while varying $I_{p}$ and $G$, respectively.
Figure 13 presents results from a trained model obtained by minimizing (28) via sparse EKI [74]. Compared to standard EKI, sparse EKI incorporates sparsity into the coefficients of the trained model via $\ell_{1}$-constrained optimization. It can be seen in Fig. 13a that the neural network captures the linear pattern of the true term $-I_{p}/t_{p}$ and achieves overall better agreement than the results in Fig. 12a. In addition, Fig. 13b shows that the neural network learns to correctly ignore $G$ in all plotted scenarios. An example of sparse EKI driving initially non-zero error model coefficients to zero is presented in Fig. 13c, for which the standard EKI would lead to larger variability with even more EKI iterations. The sparse EKI also drives some redundant coefficients of the error model to values very close to zero.
This example highlights to importance of sparsity constraints in learning error models, to avoid learning nonzero error terms that may otherwise arise because the error has zero or small projection on the output data of the model.
@@ Acknowledgments
TS, AMS, and JW were supported by Schmidt Sciences, LLC. AMS is also supported by a Department of Defense Vannevar Bush Faculty Fellowship, and by the SciAI Center, funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729. MEL was supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1745301 and by funding from the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard.
@@ Appendix A Ensemble Kalman inversion
The use of ensemble Kalman based methods for parameter calibration and the solution of inverse problems, and history of this subject, is overviewed in [Section 4][125]. To be concrete we will concentrate on a particular variant of the methodology, sometimes termed Ensemble Kalman inversion (EKI). This is a specific ensemble-based, gradient-free optimization scheme that was proposed and studied in [126]; we emphasize that other ensemble Kalman based methods share the core desirable attributes of EKI, namely that it is derivative-free, is effective with relatively few evaluations of the forward model $\mathcal{G}$ and is robust to the presence of noise in the evaluations of $\mathcal{G}.$
The core task of EKI is equivalent to a quadratic optimization problem, which facilitates adding linear equality and inequality constraints [127]. To explain the details of EKI, we first introduce a new variable $w=\mathcal{G}(\theta)$ and variables $v$ and $g(v)$:
@@
to threshold those $\theta_{i}$ with values close to zero, after having solved the constrained optimization problem in (33). Such a thresholding step after the $\ell_{1}$-constrained optimization in (33) is equivalent to adding $\ell_{0}$ constraint. More details about imposing sparsity into EKI can be found in [74].
For the multiscale Lorenz 96 example with $c=10$ and the single scale Lorenz 96 example, 100 ensembles are used and the number of EKI iterations is 20. The multiscale Lorenz 96 example with $c=3$ is more challenging to train, for which 200 ensembles are used and the number of EKI iterations is 30. The single-scale Lorenz 96 example uses 200 ensembles and the number of EKI iterations is 10. The ultradian model example is not chaotic while the convergence of standard EKI is relatively slow, for which 50 ensembles are used and the number of EKI iterations is 50. The noise level $\eta$ is estimated by running ensemble simulations of the true system and calculating the covariance of the data $y$. In practice, only the diagonal of the estimated covariance matrix is kept.
@@ Appendix B Application to a Glucose-Insulin Model
The functional forms of the parameterized processes in Eq. (27) are as follows: